A savings account earns 4.3% compounded continuously. How much would u need to deposit now in order to have a balance of 25000 after three years ? Assume that no additional deposits are to be made.
Answer by jorel1380(3719)

Continuous compounding is FV=P*e^rt, where FV equals future value, P is the amount of the initial deposit, r is the interest rate, and t is time. Let n be the amount of the initial deposit. Then:
25000=n*e^3*0.043=n*e^0.129
25000/1.1376901241657315822358920377993=n
n=21974.349 as the amount necessary for the initial deposit
☺☺☺☺